40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 43 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  494   Fri Feb 11 19:45:49 2011 FrankNotesBEATupdated diagram of current cable delay setup

cable-delay_v2.pdf

  493   Fri Feb 11 14:35:21 2011 frank, taraNotesElectronics EquipmentSR560 noise level

Disconnect the RFPD signal and terminate with a 50 ohm to measure the noise floor of the cable delay method.

Starting time : 11-2-11-22-35-33 (yy-mm-dd-hh-mm-ss)

 

  • C3:PSL-GEN_DAQ15 : SR560, DC-coupled signal, gain somewhere between 20, LP30KHz
  • C3:PSL-GEN_DAQ16 : SR560, AC-coupled signal, gain 1e4, LP30kHz

Quote:

We are measuring the noise level of SR560 at low frequency ( ~10mHz) , with the gain 1e4, AC couple, low pass 30 Hz. The  data is recorded by C3:PSL-GEN_DAQ14 channel.

 The input is terminated with a 50 ohm terminator. The setup of the SR560 under test (S/N 02785) is the same as the one we are using now.

Tomorrow we will measure the noise level of the whole setup (SR560, cables, all mixers and amplifiers included.)

 

 

  492   Fri Feb 11 13:46:20 2011 FrankNotesBEATcalibrated error signal vs frequency

plots are not taking the change of transmitted PD power when changing VCO frequency into account !

setup and cal data from post #485

 

mixer_output.png

calibrated_error_signal.png

  491   Thu Feb 10 19:37:23 2011 frank, taraNotesElectronics EquipmentSR560 noise level

We are measuring the noise level of SR560 at low frequency ( ~10mHz) , with the gain 1e4, AC couple, low pass 30 Hz. The  data is recorded by C3:PSL-GEN_DAQ14 channel.

 The input is terminated with a 50 ohm terminator. The setup of the SR560 under test (S/N 02785) is the same as the one we are using now.

Tomorrow we will measure the noise level of the whole setup (SR560, cables, all mixers and amplifiers included.)

time 11-2-11-3-32-31 (yy-mm-dd-hh-mm-ss)

 

  490   Thu Feb 10 16:45:32 2011 FrankSummaryElectronics Equipmentultra low-noise pre-amps

some old designs i've built some years ago using the cheaper DIP versions of the matched BJT pairs from Analog Devices.

Designs are DC-coupled, gain 1000 as they were designed for measuring the noise of photodetectors.
Given values are not 100% what i've used later, only for drawing the schmatics (e.g. gain setting resistors or compensation), but order of magnitude is right.

Eagle-files + Documentation:

uln-amp_NPN_v1.brd

uln-amp_NPN_v1.sch

uln-amp_PNP_v1.brd

uln-amp_PNP_v1.sch

uln-amp_NPN_v1.documentation.pdf

uln-amp_PNP_v1.documentation.pdf

BJT_preamp_inputnoise.png

  488   Thu Feb 10 12:00:55 2011 ranaNotesBEATcollection of beat data, H1 arm noise, and SR560 noise

 Here's the plot from the H2YAC doc that explains the idea.

If we can get down to ~1 Hz/rHz at 1 Hz, it would be nice for measuring the suspension performance at that point.

Attachment 1: Screen_shot_2011-02-10_at_10.39.43_AM.png
Screen_shot_2011-02-10_at_10.39.43_AM.png
  487   Thu Feb 10 01:16:42 2011 TaraNotesBEATcollection of beat data, H1 arm noise, and SR560 noise

I plot beat noise PSD, with estimated H1 arm noise, and SR560 noise converted to frequency noise in the current beat measurement. From the approximation, SR560 noise will be smaller than H1 noise.

 

 The conversion for SR560 noise to frequency noise is computed by finding the slope of signal from self beat measurement.

parameters are from Frank's entry below

amplified (DC-coupled) signal from mixer using SR560, LP@30Hz, gain20 in channel C3:PSL-GEN_DAQ15

185.0MHz :  -2.405V
154.2MHz :   0.005V
123.2MHz :   2.263V

 

The pkpk values is then (2.263V - -2.405V)/gain20 = 0.2334 V over 185 MHz - 123.2 MHz = 61.8 MHz range.

Thus the signal is = (0.2334V)/2 sin ( pi df/ 61.8 MHz)

This gives the maximum slope = 0.2334 V* pi /2 /61.8 MHz, or  ~ 170 MHz/V.

SR560 noise is estimated to be flat, 5nV, at high f, a corner at 10Hz which gives 5 uV at 10mHz.

multiplied by 17 MHz/V to get frequency noise from SR560.

 

 

Attachment 1: beat_2011_02_09.png
beat_2011_02_09.png
Attachment 2: beat_data.mat
Attachment 3: code_2011_02_09.m
load beat_data.mat

%beat_pll_100kHz(:,3)=beat_pll_100kHz(:,2).*sqrt(1+(0.16./beat_pll_100kHz(:,1)).^2);

% correction for 160mHz high pass (AC couple)
beat_pll_100kHz_2(:,3)=beat_pll_100kHz_2(:,2).*sqrt(1+(0.16./beat_pll_100kHz_2(:,1)).^2);
beat1_2011_01_30(:,3)=beat1_2011_01_30(:,2).*sqrt(1+(0.16./beat1_2011_01_30(:,1)).^2);

loglog(beat1_2011_01_30(:,1),beat1_2011_01_30(:,3)*71e3... 
      ,beat2_2010_11_29(:,1),beat2_2010_11_29(:,2)*71e3,...
... 7 more lines ...
  486   Thu Feb 10 00:21:55 2011 FrankSummaryBEATcomparison VCO feedback signal and "cable technique"

data taken at 02/06/2011 10:10:00 UTC, duration = 24h

updated plots:

cavity-drift2.png

cavities-temp.png

data files:

RCAV_TEMPAVG.mat

ACAV_TEMPAVG.mat

VCOFREQ.mat

CABLE.mat

 

structure of data files like this:

y1 =
   name: 'C3:PSL-ACAV_VCOMON'
   min: [1440x1 double]
   mean: [1440x1 double]
   max: [1440x1 double]
   rate: 0.0167
   start: 981022215
   duration: 86400
   data: [1440x1 double]

raw data in min/mean/max
calibrated (mean) data in data

  485   Wed Feb 9 19:57:43 2011 FrankNotesBEATnew cable length - calibration data

shorter cable: cable length 62 inches

amplified (DC-coupled) signal from mixer using SR560, LP@30Hz, gain20 in channel C3:PSL-GEN_DAQ15

185.0MHz :  -2.405V
154.2MHz :   0.005V
123.2MHz :   2.263V

max range of double-passed VCO signal: 142MHz-170MHz

142MHz :  1.38V
170MHz : -1.685V

As the FET preamp need some more time to set up i added the AC-coupled signal using a SR560, gain 10k, LP30Hz into C3:PSL-GEN_DAQ16

FET pre-amplified signal will be connected to C3:PSL-GEN_DAQ14   preamp broken

  484   Wed Feb 9 15:41:45 2011 frank, taraDailyProgressRefCavrunning thermal PID perl script

The thermal control perl script is added for ACAV. The file is saved in SUN machine, all channels are renamed for ACAV and the limit for hardstop is changed to 4.9 V.

an medm screen for ACAV RCPID control, C3PSL-ACAV_RCPID.adl, is also created.

Now we are tuning the gain for ACAV Temperature control.

Quote:

We installed ezca library on PLS's Sun machine. Now, we can run perl script for Refcav thermalPID control.

The PID gain is being optimized. Channels for SLOWDC's PID thermal control are also created (C3:PSL-FSS_SLOWPID....)

and saved in fss_pid.db file. 

 

We copied a command package from op440m to the sun machine here, and it can run perl script (RCthermalPID.pl)used at 40m.

The current gain setup is

KP = -0.6

KI = -0.007

KD = 0

The plot below shows a response from a step change of temperature (set to 34.86 C). The cavity temperature does not hang around the set temperature. Instead, it hovers just below the set temp point, see fig1. The time span is 110 mins.

I think it reaches equilibrium quite fast.

 

 

 

  483   Wed Feb 9 02:07:25 2011 Frank,TaraDailyProgressBEATlong-term frequency drift/noise - once more

Changed the setup to be used with two different cables at the same time, but only using one right now.

mixer signal calibration with PD signal ~4.7dBm:  5V = 157.295MHz  -5.09V=150.725MHz    dfpkpk=6.57MHz

channel names:

  • C3:PSL-GEN_DAQ15 : SR560, DC-coupled signal, gain somewhere between 100 and 200, LP30KHz
  • C3:PSL-GEN_DAQ16 : SR560, AC-coupled signal, gain 1000, LP30kHz

data valid from 11/02/09 6:20:00 UTC

will post schematic later...

 

  482   Tue Feb 8 13:50:01 2011 FrankNotesRefCavplz no more temp servo tuning - want to take data

by tuning the servo Tara unlocked both cavities and they are out of range right now, so plz no more temp servo tuning until further notice

  481   Tue Feb 8 12:02:06 2011 taraDailyProgressRefCavTemperature response

The perl script is working fine on RCAV. It has faster response than the other control system (see ACAV's temp)

Attachment 1: Screenshot.png
Screenshot.png
  480   Tue Feb 8 10:41:59 2011 FrankSummaryBEATcavity drift within 24h

i analyzed  the data taken with the short cable and compared it to the signal from the tuning input to the VCO.
I used the 9.order polynomial fit from yesterday to convert the VCOMON voltage into absolute VCO frequency.
The beat frequency is then twice that frequency.

On the other hand the cable-delay technique was sampled with channel FREQCOUNT.
I've measured the peak voltage of the amplified mixer response to be about 4.52V.
The frequency change which corresponds to peak-to-peak change in output signal (see yesterdays entry) i've used is the smallest value i measured ( 6.42MHz).

Using the sampled data, divided by 4.52V, taking sin-1 of that and multiplying it by 6.41MHz /2 gives the measured frequency change.
As i don't get an absolute frequency from that measurement i added an offset of 151.5MHz (this information is from the VCO's absolute frequency) in order to compare the fluctuations with the VCO signal.

RESULT:  beat signal fluctuations over a long period of time can be measured using the VCO feedback signal.
                Both signals absolutely agree except for the region where the mixer signal is at it's peak and so the slope is close to zero and so the uncertainty is too large.

Here the plot:

cavity-drift.png

Matlab-code:

t0 = tconvert('02/06/2011 10:10:00');
dur = 3600*24*1;

chans = {...
'C3:PSL-ACAV_VCOMON',...
'C3:PSL-FSS_FREQCOUNT',...
};

y1 = get_trend(chans{1},'minute',t0,dur);
y2 = get_trend(chans{2},'minute',t0,dur);

y1.data = y1.mean;
y2.data = y2.mean;

y1.data = 2*fittedmodel2(y1.data);
y2.data = 151.5-asin(y2.data./4.52).*3.21;

t1 = linspace(0,dur,dur*y1.rate)';
t2 = linspace(0,dur,dur*y2.rate)';

figure(1)
plot(t1,y1.data, 'r',t2,y2.data, 'b')
xlabel({'time [s]'});
ylabel({'beat frequency [MHz]'});
grid

  479   Tue Feb 8 00:26:44 2011 frank, taraDailyProgressRefCavrunning thermal PID perl script

We installed ezca library on PLS's Sun machine. Now, we can run perl script for Refcav thermalPID control.

The PID gain is being optimized. Channels for SLOWDC's PID thermal control are also created (C3:PSL-FSS_SLOWPID....)

and saved in fss_pid.db file. 

 

We copied a command package from op440m to the sun machine here, and it can run perl script (RCthermalPID.pl)used at 40m.

The current gain setup is

KP = -0.6

KI = -0.007

KD = 0

The plot below shows a response from a step change of temperature (set to 34.86 C). The cavity temperature does not hang around the set temperature. Instead, it hovers just below the set temp point, see fig1. The time span is 110 mins.

I think it reaches equilibrium quite fast.

 

 

Attachment 1: 2011_02_07_RCPID.png
2011_02_07_RCPID.png
  478   Tue Feb 8 00:03:02 2011 FrankSummaryVCOcalibrated VCO monitor signal now fully available

the configuration is currently running with a softIOC in /caltech/target/SoftIOC/ on fb2, the database is PSL.db.
Will move that on the VME crate when we reboot the crate the next time.

Currently two channels are defined: C3:PSL-ACAV_VCOMON_CAL and C3:PSL-ACAV_VCOFREQ

C3:PSL-ACAV_VCOMON_CAL contains the polynomial fit for the VCO monitor voltage, which is good for -4.3V to 5.0V in tuning voltage.
To extend the range down to -5V i've used a second calc function named
C3:PSL-ACAV_VCOFREQ.
The frequency value where the polynomial function crosses the measured function of the VCO tuning is about 71.6MHz.
As the tuning range ends here the change in frequency for even smaller tuning voltages is almost zero.
So the second function simply checks if the frequency is greater. If so, the output is the polynom. If smaller the value is fixed to 71.6MHz.

Both channels are written to the frames from now on.

 

PSL.db:

record(calc,"C3:PSL-ACAV_VCOMON_CAL")
{
        field(SCAN,".1 second")
        field(DESC,"VCO frequency")
        field(EGU,"MHz")
        field(PREC,"4")
        field(HOPR,"90")
        field(LOPR,"70")
        field(INPA,"C3:PSL-ACAV_VCOMON.VAL")
        field(INPF,"0.0005513")
        field(INPG,"-0.003731")
        field(INPH,"-0.005932")
        field(INPI,"-0.02587")
        field(INPJ,"1.406")
        field(INPK,"79.99")
        field(CALC,"F*A^5+G*A^4+H*A^3+I*A^2+J*A+K")
}

record(calc,"C3:PSL-ACAV_VCOFREQ")
{
        field(SCAN,".1 second")
        field(DESC,"VCO frequency")
        field(EGU,"MHz")
        field(PREC,"4")
        field(HOPR,"90")
        field(LOPR,"70")
        field(INPA,"C3:PSL-ACAV_VCOMON_CAL.VAL")
        field(INPB,"71.6")
        field(CALC," A>B?A:B")
}

  477   Mon Feb 7 20:33:00 2011 FrankSummaryVCOcalibrated VCO monitor signal now available

channel name is C3:PSL-ACAV_VCOMON_CAL, calibrated in MHz using the fifth order polynomial equation at the moment.
Unfortunately the field length for the EPICS CALC record is limited (but not mentioned anywhere in the manual !) so the 9th order equation is too long

will change that later to 2 channels for calculating parts of the equation each and then finally combining those parts in a third record to C3:PSL-ACAV_VCOFREQ

  476   Mon Feb 7 18:13:31 2011 FrankSummaryVCOVCO tuning - fitted function

tried to fit  the tuning of the VCO in order to calibrate the VCO tuning voltage into frequency shift.
I've only fitted it down to -4.3V using 5 degrees as there is no change in frequency below that point anymore.
Will add an EPICS software channel which contains the calibrated data.

Linear model Poly5:
     f(x) = p1*x^5 + p2*x^4 + p3*x^3 + p4*x^2 + p5*x + p6
Coefficients:
       p1 =   0.0005513
       p2 =   -0.003731
       p3 =   -0.005932
       p4 =    -0.02587
       p5 =       1.406
       p6 =       79.99

vco-tuning_fit1.png

Using 9 degrees the fit is not as good in the valid region for the 5 degrees, but covers the entire range from -5V to 5V

Linear model Poly9:
     f(x) = p1*x^9 + p2*x^8 + p3*x^7 + p4*x^6 +
                    p5*x^5 + p6*x^4 + p7*x^3 + p8*x^2 + p9*x + p10
Coefficients:
       p1 =   4.16e-006
       p2 = -1.095e-006
       p3 =  -0.0003138
       p4 =   0.0003221
       p5 =    0.007185
       p6 =    -0.01047
       p7 =    -0.05412
       p8 =    0.006482
       p9 =       1.498
       p10 =       79.97

vco-tuning_fit2.png

  475   Mon Feb 7 16:14:19 2011 FrankNotesBEATcable phase noise measurement calibration

used two frequency ranges to calibrate the mixer signal

  1. calibration factor is 6.85MHzpkpk measured between 150MHz and 160MHz
  2. calibration factor is 6.42MHzpkpk measured between 147MHz and 154MHz

those numbers should be sufficient for comparison with the VCO feedback signal.

  474   Mon Feb 7 15:38:38 2011 FrankNotesBEATcable phase noise measurement stopped

stopped taking data for data calibration at 11/02/07 23:35:00 UTC

  473   Sun Feb 6 02:38:07 2011 FrankDailyProgressRefCavboth cavities locked again

locked both cavities. Current slowdc value is 0.1925. VCOmon is -2.2V. Changed RCtemp to 35.00.

Data is valid since 11/2/6 10:10:00 UTC

  472   Fri Feb 4 19:26:50 2011 FrankDailyProgressBEATchanged to shorter cable

exchanged the 500ft spool by a shorter cable to get more range (but less resolution).
Due to the lower losses of the cable i also removed the 2W amplifier.
Right now both cavities can't be locked at the same time, they are slightly out of range.

RCAV is resonant at 0.1958 for the slow actuator, ACAV is resonant at 0.1930

max range for VCO is reached at 0.1948

i've changed the RCAV settemp a little bit and will keep an eye on that and we will hopefully be back online tonight

  471   Fri Feb 4 13:59:48 2011 FrankDailyProgressBEATtrend of the setup

 This setup has too much gain (i.e. not enough range). Please reduce the arm length asymmetry by a factor of 10 so that we can monitor over 24 hours.

Also the temperature channels ought to be calibrated (via the EPICS .db) so that the readout is in degC instead of ARB.

Attachment 1: Untitled.png
Untitled.png
  470   Fri Feb 4 01:03:47 2011 FrankDailyProgressBEATcable delay setup

changed the frequency noise readout to the cable-delay version to see how it works.
As the loss of the cable is very large and the signal from the photodetector not very strong i had to add some amplifiers for signal conditioning (see figure below).
Data is acquired with channel C3:PSL-FSS_FREQCOUNT in replacement for the frequency counter. Started taking data around 11/2/4 4:16:30 UTC.
Will do final calibration tomorrow. First test gave signal from mixer changes from peak-to-peak for about 600kHz (+/-10%) in frequency change.

cable-delay.png

  469   Thu Feb 3 14:25:18 2011 FrankSummaryBEATFrequency counter time series

2 stretches of data taken at the following UTC times:

11/2/2 23:18:19  duration: ~9min

11/2/2 23:22:50  duration: ~9min

units for data: volts  SCALE: 10kHz/V

 

Here a screenshot for the first stretch from dataviewer:

frequcount_timeseries2.png

 here the data: freqcounter_timeseries.dat

 we didn't take longer timeseries with more range as the resolution is so bad that it couldn't be used to extract a spectrum down to mHz .

RAW data can be accessed via 131.215.114.84:8088 or SSH-login and doing the usual.
Channel name is C3:PSL-FSS_FREQCOUNTER

 

 

  468   Wed Feb 2 20:47:00 2011 koji, frank, taraNotesNoiseBudgetmechanical peaks in beat noise

We were looking for optics which cause peaks in beat noise. Periscopes and mirror infront of RCAV are more sensitive

than other components. Also, cavity's suspension (pendulum) might be needed to be modified.

 

Mechanical peaks from optical components can be up/down converted and cause extra noise in 

beat measurement, thus we also need to eliminate them. By listening to the PLL feedback signal, we

can tap each optic component to verify which one is sensitive to acoustic noise. So far there are

periscopes, mirror in front of RCAV, and cavity suspension that are quite sensitive to acoustic noise

compared to the others. For the suspension, we might have to mount the cavity on a block instead of

hang it on springs.

 

After we worked around the table, RCAV temp jumped again. By kicking the table, we can bring RCAV temp

back to what it should be (fig1 top left panel.) ACAV and RCAV RCTRANSPD also changed as well (two right panels.) 

This might be electronics jump, or that the servo was oscillating before, and when the

kick brought it back to lock again. At least ACAV temp read out is quite immune to these jumps. 

Attachment 1: 2011_02_02_kick.png
2011_02_02_kick.png
  467   Wed Feb 2 17:59:24 2011 ranaNotesNoiseBudgetTemperature vs frequency shift

 That's nice, but let's see the time series of the frequency counter / PFD data as well as the temperature and VCOmons.

  466   Wed Feb 2 14:18:37 2011 taraNotesNoiseBudgetTemperature vs frequency shift

 I plotted the temperature difference between two  cavities vs VCOMON.

This might give us how temperature on the vacuum can couples to frequency drift (at least for diurnal change.)

More data is needed to verify this.

 

Since VCOMON channel was added yesterday, I can plotted the data upto 20hrs.

 

Fig1 shows time series data of 1) temperature difference between ACAV and RCAV , 2)VCOMON.

the temperature is multiplied by a factor of 100 for easy comparison.

 

Fig2 is a plot of dT vs VCOMON. The data before t = 500 min are discarded (no VCOMON data yet).

More data is needed for a better linear fit.

 

Electronic jumps in RCAV temp results in the jumps in dT as well. We can see jumps at t ~300, 700, 800 in fig1.

We'll figure out how to fix this jump.

Attachment 1: temp-vco.png
temp-vco.png
Attachment 2: vco_vs_T.png
vco_vs_T.png
  465   Wed Feb 2 13:48:22 2011 taraDailyProgressBEATComparison bewteen beat from PLL and FC

I used PLL feedback to measure beat noise from dc to 3 Hz and dc to 1.5 Hz to compare with the data taken yesterday by FC.

 

There is no peak around 3 Hz (probably a peak from cavity suspension)in FC data, I'm not quite sure if we what we see is real or not.

 

Attachment 1: beat_2011_02_02.png
beat_2011_02_02.png
  464   Tue Feb 1 23:16:10 2011 FrankHowToElectronics EquipmentTF (attenuation) of 500ft of RG58

I've measured the TF of 500ft of RG58C/U cable to see if the loss is about the same calculated by that piece of software i've found.
I've measured 33.4dB attenuation for the LCOM cable, the calculated value is 32.1dB for some unknown RG58C/U cable.
So i think we should go ahead and calculate the required length and try it.

TF_500ft_RG58CU.png

This is a picture of the 500ft spool we have. As you can see there is only about 1.5inch on that spool, the rest is empty. Height is a few inches.
So regarding the size we can easily have several 100m of cable in a small package

P1700783.JPG

  463   Tue Feb 1 22:22:46 2011 frank, taraNotesBEATbeat signal down to 12.5 mHz

We measured beat noise from the frequency counter(FC) instead of the feedback from PLL. The result is plotted below

 

We switch to use the FC because we hope it will allow us to measure the signal down to lower frequency (~10 mHz) which is not quite possible

for PLL because of its small input frequency range for the acceptable phase noise level.

 

The FC can measure the relative frequency from the chosen center frequency and give a voltage output (from 0 to 8 V.)

For example if we set the center frequency to be 160 MHz, with 1MHz/V gain, signals at 156 MHz and 164 MHz will correspond to 0 V and 8 V respectively.

Since the LIGO VCO range is ~ 10 MHz, we tried measuring the signal with 500 kHz/V setup which is equal to 4 MHz range. However the noise is too high,

so we have to choose a new gain setup, and  20kHz/V setup is still acceptable for our signal compared to the signal from PLL.

 (It turns out that data from 50kHz/V is too noisy)

Attachment 1:

Blue and Green show the beat noise measurements from PLL and FC. They agree well. The gain for FC is set to 2kHz/V.

When the temperature became more stable, the data could be acquired to down to ~ 3mHz (red).

 

We used AC coupling for FFT measurement, so the TF for AC coupling is measured. There is a 160 mHZ high pass, and the data is corrected accordingly.

Attachment2:

Time series for ACAV's temperature, RCAV's temperature, VCO mon, and Frequency count Vout during the FFT measurement.

 

Attachment3: show TF measurement for AC highpass and fit with 160mHz high pass.

The plot is flipped because chA is AC couple and ChB is DC couple during the measurement, and

The output is B/A.

Quote:

This morning I was able to measure the beat noise down to 12.5 mHz. So I plot it together with the noise budget here.

We also added channel for frequency counter,C3:PSL-FSS_FREQCOUNT which will allows longer data acquisition time for lower frequency.

Then we will be able to see the temperature effect at lower frequency(~ 10 mHz.)

 

Fig1: beat signal at 12.5 mHz to 10 Hz

 

We are also working on PID thermal control for refcav.

Perl script for PID thermal control won't work on Solaris because it doesnt have ezcaread/write command, we will get that from op40m machine at the 40m.

(We can't run Perl script on Linux because it complains that it fails to read / write data from C3:PSL-FSS_HEATER.)

 

 

Attachment 1: PLL_FC_good.png
PLL_FC_good.png
Attachment 2: 2011_02_01_log3.png
2011_02_01_log3.png
Attachment 3: ac_highpass.png
ac_highpass.png
Attachment 4: 2011_02_01_dat.mat
Attachment 5: code_2011_02_01.m
load 2011_02_01_dat.mat

fc=fc_20kHz_22db_gd;
%correct for 160 mHz high pass
fc(:,3)=fc(:,2).* sqrt( (1 + (0.16./fc(:,1)).^2 ));

loglog(fc_2kHzv(:,1),fc_2kHzv(:,2)*2e3,...
        beat5Hz(:,1),beat5Hz(:,2)*71e3,...
        t2kHzv(:,1),t2kHzv(:,2)*2e3,...
... 5 more lines ...
  462   Tue Feb 1 20:47:14 2011 FrankNotesDAQVCO feedback signal now recording

added the VCO feedback signal to the FB. We had it connected but not written into the frames so far

Channel name is C3:PSL-ACAV_VCOMON

  461   Tue Feb 1 19:26:47 2011 FrankNotesBEATloss of RF cables

found a nice calculator here:

http://vk1od.net/calc/tl/tllc.php

which has a huge amount of different cable types in it's database. Checked/compared the calculated values for some examples given in several datasheets found on the web and they are close within each other.
Here an example for our current case for the 500ft spool of RG58C/U we have in the lab. Loss is only 32dB which looks pretty good to me for that cheap cable.

Parameters  
Transmission Line RG-58C/U
Code RG-58C/U
Data source DSE
Frequency 160.000 MHz
Length 500.000 ft
Results  
Zo 50.00-j0.13 Ω
Velocity Factor, VF -2 0.660, 2.293
Length 44334.70 °, 123.152 λ, 152.400 m
Line Loss (matched) 32.146 dB
Loss model source data frequency range 50.000 MHz - 1000.000 MHz
Correlation coefficient (r) 0.999926
  460   Mon Jan 31 21:40:50 2011 Frank, TaraSummaryDAQnew channel for frequency counter added

We added a new channel C3:PSL-FSS_FREQCOUNT which contains the analog output of the SR620 frequency counter connected to the beat photodiode.

 

  • C3:PSL-FSS_FREQCOUNT is connected to CH6 of the first 3123 card, connector P3 (first D-SUB). This is also J4 of the PSL x-connect.
  • Added channel entry in fss.db database. Sampling rate is 10Hz.  Scaled from -10 to 10V in unit  "volts" as we don't know which settings we will use later.
  • Also added channel to daqd. restarted daqd. Manual restart of daqd as controls using script requires password for controls.
    That shouldn't be so but i was to lazy to search for the bur

 

Counter output is 0 to 8V, right now used in relative mode which means zero frequency deviation equals 4V. Current gain for first test is 100kHz/V.

  459   Mon Jan 31 20:24:45 2011 frank, taraNotesBEATbeat signal down to 12.5 mHz

This morning I was able to measure the beat noise down to 12.5 mHz. So I plot it together with the noise budget here.

We also added channel for frequency counter,C3:PSL-FSS_FREQCOUNT which will allows longer data acquisition time for lower frequency.

Then we will be able to see the temperature effect at lower frequency(~ 10 mHz.)

 

Fig1: beat signal at 12.5 mHz to 10 Hz

 

We are also working on PID thermal control for refcav.

Perl script for PID thermal control won't work on Solaris because it doesnt have ezcaread/write command, we will get that from op40m machine at the 40m.

(We can't run Perl script on Linux because it complains that it fails to read / write data from C3:PSL-FSS_HEATER.)

 

Attachment 1: beat_2011_01_30_lowf.png
beat_2011_01_30_lowf.png
  458   Mon Jan 31 19:42:36 2011 FrankNotesComputerscurrent IP-address

for fb2 from outside is

131.215.114.84

  457   Sun Jan 30 22:58:35 2011 TaraDailyProgressElectronics EquipmentRCAV RFPD fixed

I adjusted the temperature on ACAV and RCAV so that both cavities can be locked simultaneously.

I measured the beat noise and see improvement at low frequency after we fixed the RFPD.

 

After we reinstalling the RFPDs on the table, we need to wait for the temperature to settle before both cavities can be locked.

(The temperature readout jumps by ~0.05 degree if we work on the table. The cause is not known yet. we tried pulling the cables for T readout,

but this does nothing to the T readout.  This problem will be investigated soon.)

I check the beat noise and see some improvements at low frequency . It does not look a lot for now, but I think it might go lower

if I re align the beam to the cavities and the beam for beat measurement. It might drift away a bit during the break, and I haven't checked that yet.

 

The input power is ~ 1 mW for both cavities.

RCAV gain = 25

ACAV gain= 8.0 on the knob.

 

 

Quote:

The TF of 2 35.5MHz RFPD and 21.5 MHz RFPD are measured by the Jenne laser. RCAV's RFPD has the better response, while ACAV's RFPD peak is slightly shift to 37 MHz. But the peak is low Q so we lose only a few dB. PMC's RFPD is now tune to have a resonance peak at 21.5 MHz and a notch at 43 MHz as it should be.

 

 

Attachment 1: beat_2011_01_30.png
beat_2011_01_30.png
  456   Fri Jan 28 13:35:35 2011 Tara, FrankDailyProgressElectronics EquipmentRCAV RFPD fixed

The TF of 2 35.5MHz RFPD and 21.5 MHz RFPD are measured by the Jenne laser. RCAV's RFPD has the better response, while ACAV's RFPD peak is slightly shift to 37 MHz. But the peak is low Q so we lose only a few dB. PMC's RFPD is now tune to have a resonance peak at 21.5 MHz and a notch at 43 MHz as it should be.

 

Attachment 1: 35MHzPD.png
35MHzPD.png
Attachment 2: pd21.png
pd21.png
  455   Wed Jan 26 20:21:57 2011 FrankSummaryElectronics EquipmentACAV RFPD modified

The ACAV RFPD stopped working this afternoon. It had high current consumption on the +15V supply, causing the supply to drop down to 2.3V.
I turned out that the logic IC (U8, see schematic) was broken and so the +5V (internally regulated) caused the high current flow which is supplied from the +15V.
I removed U8 entirely as we don't need it. Pin7 of U2 can be left open according to the datasheet in order to enable the device. Diode is now working again.

Already yesterday we made several changes in order to make the RFPDs for both cavities the same. Basically more DC and AC gain.
Attached the modified schematic (only page 1) for ACAV RFPD. Changes are in red.

D980454-00_modified_ACAV.pdf

 

  454   Tue Jan 25 16:30:16 2011 Tara, FrankDailyProgressElectronics EquipmentRCAV RFPD fixed

we replaced the photodiode in the RFPF fro the reference cavity (RCAV). The old one looked like shit. Below pictures of the old and new photodiode.

RCAV-RFPD1.jpg

 old photodiode - picture 1

 

RCAV-RFPD2.jpg

old photodiode - closeup view

 

new_RCAV-RFPD3.jpg

new photodiode

  453   Tue Jan 25 11:30:05 2011 FrankNotesElectronics EquipmentPerkin Elmer 2mm InGaAs photodiodes - dark current characterization

PerkinElmer_2mm.png

  452   Mon Jan 24 21:02:15 2011 taraDailyProgressNoiseBudgetRIN coupling to frequency noise

I'm measuring the TF between ACAV (ref) and  feedback to VCO (response).

The TF looks ok, but I still need to verify by increasing the laser power and see if the TF goes up with the same factor.

 

We are determining the coupling coefficient from intensity to frequency change via thermo-optic effect.

Using a swept sine measurement with 4000 integration cycles, from 10Hz to 100kHz, to measure the TF 

between A) ACAV RCTRANSPD (ref) and B) feedback to VCO(response), we can see some reasonable results.

I tried 2000 integration cycles with feedback to PZT, but it seems the laser is too noisy. I might try longer integration time later.

 

To make sure that the measured TF is real, we must check that when the power to the cavity is increased, the magnitude of

the TF should increase by the same factor. 

Once we have the TF between ACAV_RCTRANSPD and VCO feedback, we can calculate how RIN would couple in to frequency noise by

Thermo optic noise (RIN) [Hz] = VCO calibration [Hz/V] x TF between VCO feedback and ACAV_RCTRANSPD [V/V] x

                                                    VCO TF [V/V]    x RIN x DC level of ACAV_RCTRANSPD

 

Note the TF of the VCO box (VCO TF) is measured and fit. One pole is at 1.25 Hz, and one zero is at 38 Hz. see fig 1.

 VCO's calibration for V to Hz can be found here, The temperature is quite stable and VCO_MON has been around -0.7 to -0.3 V, we can approximate to be 1.4 MHz/V for now.

Attachment 1: vco_tf_fit.png
vco_tf_fit.png
  451   Mon Jan 24 20:26:44 2011 taraNotesElectronics EquipmentRFPD photodiodes

I also got the Jenne laser and its power supply from 40m for TF measurement. They are on the PSL table.

Quote:

got some 2mm InGaAs photodiodes from Peter. So we can go ahead and replace the dirty one from the ACAV RFPD tomorrow and re-tune it to 35.5MHz.

 

  450   Mon Jan 24 19:05:41 2011 FrankNotesElectronics EquipmentRFPD photodiodes

got some 2mm InGaAs photodiodes from Peter. So we can go ahead and replace the dirty one from the ACAV RFPD tomorrow and re-tune it to 35.5MHz.

  449   Thu Jan 20 11:32:40 2011 taraPhotosBEATbeat layout with 3" height

The picture is posted on the board outside the lab as well.

Attachment 1: beat.jpg
beat.jpg
  448   Mon Jan 10 14:27:11 2011 JanSummaryPEMWeekly plots; starting date 01/01/2011

 So I don't know what's going on, but the table is doing something really terrible. Maybe it vibrates as a response to ventilation, whereas the coupling of ventilation to ground is comparatively weak. Anyway, this is the end of PSL seismic measurements. Data will be kept on one of the machines in 058f. In principle, I could convert the data into frames and copy it to another location, but it may be too much work. I can always easily produce .mat files with averaged spectra etc. Let me know.

Plot7.jpgPlot9.jpg

  447   Mon Jan 10 13:28:36 2011 JanSummaryPEMWeekly plots; start date 2010-12-25

These two plots are just for completeness. In a few minutes I will be able to post the plots for the third week, which could be more interesting since it contains the transition from table to ground. I don't understand why the low-frequency disturbance is "out of phase" with the high-frequency one. It's maybe the last sign of Tara before he left, so he seems to be out of synch with the rest of the world.

Plot10.jpgPlot12.jpg

  446   Mon Jan 10 10:37:43 2011 JanNotesPEMTable versus ground

I have taken a few days of data with the T240 standing on the ground (and not on the optical table). I expected horizontal particle trajectories to lose some of their ellipticity. And that's indeed the case (left ground, right table):

Plot16_2011-01-08.jpgPlot16_2010-12-18.jpg

So the table is doing something to the displacement at high frequencies. Note that you cannot compare absolute displacements in these plots. They are individually normalized so that the maximum dispalcement during that time seen in any of the channels is equal to one. You can also compare H/V ratios (left ground, right optical table):

Plot5_2011-01-08.jpgPlot5_2010-12-19.jpg

The 0.1Hz-1Hz peak in the table data is gone. My first thought was that this may be related to the table, but the peak height also varied during the table measurement and is getting stronger again right now. So maybe there is a table effect, probably not (too low frequencies). Moving the seismometer to the ground, the H/V ratio above 1Hz has dropped to something significantly below 1. In principle you can use this information to measure the Poisson ratio of the ground, but only by making a few assumptions: (a) the spectrum is determined by Rayleigh waves, (b) the ground is half-homogeneous. In this case, the dashed line in the following figure tells us what the Poissson ratio is:

Rayleigh_HV.png

The H/V ratio should be frequency independent if the two assumptions were true. This is not the case, so our assumptions are wrong. Only the 1Hz to 10Hz range seems ok. Below 0.1Hz we sort of know that the seismic displacement is not driven by Rayleigh waves, but local sources causing heavy tilts that cannot be associated with Rayleigh waves (which is quite a non-trivial statement). Maybe we should also not forget that it may be possible that temperature changes, air currents and pressure changes may also couple into the low-frequency data. This is always the problem at the surface. You need to spend a lot of money to make good seismic measurements above ground. Between 0.1Hz and 1Hz it is also not completely surprising that the model fails (the ratio is too close to 1) since large scale heterogeneities break the half-homogeneity of the model. The weird structure above 10Hz is mysterious, but again since seismicity is likely driven by local sources at these frequencies, one would not expect a simple picture to form.

  445   Tue Dec 28 05:24:16 2010 ranaDailyProgressNoiseBudgetRIN coupling to frequency noise

To get a sensitive measurement, you should measure a coherent transfer function, not just noise. We expect the thermo-elastic effect to be linear. Try making a swept-sine measurement and fiddle with the integration time to get a measurement. Even it ends up being a very small signal, you can use the measurement parameters and the HP Application Note #243 to set an upper limit on the coupling.

  444   Mon Dec 27 23:11:32 2010 taraDailyProgressNoiseBudgetRIN coupling to frequency noise

I modulated laser power to see how it causes frequency noise in the cavity, no sign of RIN induced frequency noise has been seen yet.

 

We have been thinking about how RIN will cause random absorption on the mirror surface and turn into thermo elastic/ thermo refractive noise in the cavity.

So we try to observe it, by modulating the laser power with an EAOM. I used white noise, generated by SR785 at 2.5V, as a source to modulate the intensity.

  The power input to both cavities are 3mW. This corresponds to ~15mW total power before the BS (3 to RCAV, 12 to ACAV path).

 

1)The intensity changes as we can see from ACAV_trans_PD (fig1).

The lower curve is the power fluctuation of the laser without the modulation. The upper curve is the power fluctuation with the added white noise.

So we are sure that intensity does change from our modulation.

 

2)Then I measured the feedback signal to PZT and to AOM, with and without external white noise.  These signals represent the frequency change of the laser.

However no change observed in both signals. I plotted only the feedback signal to VCO(to AOM), since this signal should be more sensitive to change than feedback to PZT, as the

beam to ACAV path is stabilized to the RCAV already.

 

3) There is no observed effect from RIN on beat measurement yet, see fig 3.

 

The laser power output is ~54 mW, so the maximum power we can get to each cavity equally is ~10 mW.

I will try to increase more power to the cavity, as the effect should be proportional to the input power

Actually, we don't need both cavities to have the same power in order to measure the feedback signal. I can try

1) lock the laser to RCAV only with power of 10mW or upto 40mW and measure the feedback signal to PZT

2) lock both cavities, have maximum power to ACAV (could be upto 12,13 mW, the efficiency is ~25%) and measure the feedback to VCO

Attachment 1: ACAV_RIN.pdf
ACAV_RIN.pdf
Attachment 2: fdbk_to_VCO.pdf
fdbk_to_VCO.pdf
Attachment 3: beat.pdf
beat.pdf
ELOG V3.1.3-